How many grams of CO are needed to react with 3.32 g of Fe2O3? Express the mass in grams to three significant figures.


1
Expert's answer
2022-02-10T18:35:03-0500

Here is the balanced equation:

Fe2O3 + 3CO --> 2Fe + 3CO2

The molar masses are:

Fe2O3 - 159.69 g/mol;

CO - 28.01 g/mol.


"m(CO)=3.32\ g(Fe_2O_3)\times\frac{1\ mol(Fe_2O_3)}{159.69\ g(Fe_2O_3)}\times\frac{3\ mol(CO)}{1\ mol(Fe_2O_3)}\times\frac{28.01\ g(CO)}{1\ mol(CO)}=1.75\ g"


Answer: 1.75 g